# Developing Safe and Responsible Large Language Models - A Comprehensive Framework

### v1

@@ Abstract
Given the growing concerns around the safety and risks of Large Language Models (LLMs), it is essential to develop methods for mitigating these issues. We introduce Safe and Responsible Large Language Model (SR) , a model designed to enhance the safety of language generation using LLMs. Our approach incorporates a comprehensive LLM safety risk taxonomy and utilizes a dataset annotated by experts that align with this taxonomy. SR is designed to identify potentially unsafe content and produce benign variations. It employs instruction-based and parameter-efficient fine-tuning methods, making the model not only effective in enhancing safety but also resource-efficient and straightforward to adjust. Through our testing on five benchmark datasets and two proprietary datasets, we observed notable reductions in the generation of unsafe content. Moreover, following the implementation of safety measures, there was a significant improvement in the production of safe content. We detail our fine-tuning processes and how we benchmark safety for SR with the community engagement and promote the responsible advancement of LLMs. All the data and code are available anonymous at https://github.com/shainarazavi/Safe-Responsible-LLM .
keywords:
Large Language model , Responsible AI , AI Safety , Generative AI[inst1]organization=Vector Institute for Artificial Intelligence,city=Toronto, postcode=M5G 1M1, state=Ontario, country=Canada
[inst2]organization=University of Limerick,city=Castletroy, postcode=V94 T9PX, state=Limerick, country=Ireland
@@ 1 Introduction
The rise of generative artificial intelligence (AI) models has raised concerns about their potential risks to produce undesirable content due to misalignment with human values [bender_dangers_2021]. These AI risks fall into two categories: established risks, encompassing social and ethical issues like biases and misinformation, and anticipated risks, including autonomy and deceptive behaviors [weidinger_ethical_2021]. Large Language Model (LLM) alignment ensures models act ethically and produce unbiased, non-toxic content, aligned with human values [guo_evaluating_2023].
[…]
@@
LLM safety involves strategies to make LLMs safer by ensuring they generate ethical and safe language while reducing biases, harmful content, and unintended effects [touvron_llama_2023, inan_llama_2023]. Enhancing the safety of LLMs requires a comprehensive and multifaceted approach to address the generation of harmful, biased, or misleading content. Initial strategies for LLM safety involve the use of guardrails [guardrails_ai_guardrails_2024], content moderation, and safety-specific data instructions [wang2024decodingtrust] to guide the pre-training process and mitigate model bias. Techniques such as red-teaming [ganguli_red_2022], pre-training with human feedback [korbak_pretraining_2023], and data augmentation [prabhumoye_adding_2023] are employed to further minimize risks.
During the fine-tuning phase, advanced methods like instruction tuning [ouyang_training_2022, qi_fine-tuning_2023], reinforcement learning from human feedback (RLHF) [bai_training_2022, ouyang_training_2022], and safety context distillation [qi_fine-tuning_2023] (these methods are used in LLama2 safety fine-tuning too), are utilized to curb unsafe behaviors, including biases and toxicity. These methods not only ensure safer operation but also enhance the generalization capabilities of LLMs, improving their adaptability and efficiency through few-shot learning abilities [brown_language_2020]. Nevertheless, recent research in AI safety has led to the development of further enhanced training methods for LLMs. These methods include adversarial demonstrations [zou_universal_2023] to guard against attacks, the creation of counterfactual data [chen_disco_2023, fryer_flexible_2022] to avoid unreliable patterns, and the use of bias-neutralizing prompts [berg_prompt_2022] alongside adversarial training to reduce social biases [qi2023visual].
Despite these efforts, some recent studies [wang2024decodingtrust, zou_universal_2023] reveal vulnerabilities in LLMs, such as LLama 2 and GPT-4, which can be disrupted by a small number of adverse examples. The practice of prompt injection [zou_universal_2023], where specific inputs are inserted to manipulate LLM outputs, presents additional safety challenges, including privacy breaches and the spread of misinformation. Recently, jailbreak prompts[liu_jailbreaking_2023, qi_fine-tuning_2023, liu_prompt_2023] that aim to bypass language model restrictions and potentially prompting unethical outputs are the concerns, even in the most safe models including GPT-3.5/4, Vicuna, and PaLM-2 [chao2023jailbreaking]. This highlights a research gap and ongoing challenges in the safety-focused fine-tuning process for LLMs.
Our study contributes to seminal AI safety research by emphasizing the critical role of preparing specific safety-focused datasets for instruction-tuning of LLMs. To the best of our knowledge, there are no existing datasets tailored specifically for the safety-focused fine-tuning of LLMs. In particular, we have introduced a dataset rich in safety-oriented examples for this tuning phase, aiming to specifically address and mitigate prevalent LLM safety concerns. Through fine-tuning LLMs with a balanced mix of instruction data, including demonstration of both unsafe and their benign (safe) variations, our findings show the model improved capability in correcting the inherent biases. The inspiration behind our methodology comes from the principles of counterfactual data generation [madaan_generate_2021], where generating data samples that represent alternative scenarios enhances a model’s adaptability and understanding.
@@
Development of a Safety Risks Taxonomy aimed at systematically identifying and categorizing risks in LLM outputs into unsafe categories, such as bias, toxicity, stereotyping, and harm 111Important terms utilized throughout this work are given in the Glossary (see G).
@@
Introducing SR , a safety-tuned LLM on the top of the Llama2-7B-Chat [touvron_llama_2023] model that is focused on improving safety and ethical alignment through instruction fine-tuning. This approach distinguishes itself from other models, like Llama-guard [inan_llama_2023] that use a safety taxonomy primarily for classification, and [bianchi2023safety] that include a limited set of safety-focused demonstrations during instruction-tuning. In contrast, SR incorporates our safety taxonomy into the fine-tuning process. This enables SR to effectively identify and modify unsafe text into benign variations as per different types of LLM risks, while maintaining high model performance and understanding.
4.
@@
The rest of the paper is organized as: Section 2 presents the framework, Section 3 discusses related work, Section 4 outlines the experimental setup, Section 5 shows the results, , and Section 6 concludes.
2 Related Work
The exploration of safety in language models is multifaceted, focusing on methodologies ranging from embedding space adjustments to data augmentation and fine-tuning techniques. Embedding space adjustments often involve post-processing techniques to modify sentence representations and word embeddings, aiming to mitigate inherent biases [liang_towards_2020, ungless_robust_2022, ethayarajh_understanding_2019, may_measuring_2019, zhao_learning_2018]. This includes efforts to create unbiased linguistic spaces through subtraction-based methods [bolukbasi_man_2016]. In parallel, data augmentation strategies, such as the replacement of gendered words, highlight the necessity of model re-training to accommodate non-biased language [zhao_gender_2019]. Additionally, fine-tuning practices have evolved to incorporate safety interventions more efficiently, showing the community growing emphasis on minimizing bias through model adjustments [joniak_gender_2022, gira_debiasing_2022, ranaldi_trip_2023].
Prompt-based methods for bias detection have also gained attention, with research exploring bias identification through instruct-based approaches, demonstrating their effectiveness in pinpointing and addressing biases within language models [fu_gptscore_2023, inan_llama_2023, askell2021general, berg_prompt_2022, zou_universal_2023, morabito_debiasing_2023, schick_self-diagnosis_2021]. This highlights a trend towards interactive and dynamic bias mitigation strategies, which rely on real-time model responses to specifically crafted prompts or instructions.
[…]
Distinct from prior research, we employed instruction fine-tuning on an LLM using a custom-designed dataset specifically curated for enhancing safety. Whereas previous works have used small demonstrations [wang2024decodingtrust, bianchi2023safety] to guide the model toward safety, in contrast, we instruction fine-tuned for reducing bias and toxicity for safe content generation. This approach is a contribution to the field, diverging from traditional methods by focusing on targeted instruction-based fine-tuning as a means to directly incorporate safety protocols within the LLM framework.
3 Safe and Responsible Large Language Model (SR) - Framework for Ethical and Safe Large Language Models
The SR framework integrates ethical considerations and safety measures into the instruction fine-tuning of LLMs. As shown in Figure 1, the framework includes a ‘Safety Risk Taxonomy’ for categorizing specific safety risks in LLM outputs, CMD- a dataset for instruction fine-tuning, and the SR model, which is a LLM instruction fine-tuned on this dataset. The goal of SR is to align LLMs with human values for enhanced user safety.
@@ 3.1 Safety Risks Taxonomy
We define LLM Safety Risks Taxonomy based on extensive literature [qi_fine-tuning_2023, wolf_fundamental_2023, inan_llama_2023, touvron_llama_2023] outlining crucial areas of concern regarding the safe deployment and use of LLMs. These LLM key issues can arise during development and deployment and are defined as: Bias is a significant concern, where the goal is to ensure LLM produce content that is balanced and fair, avoiding favoritism towards any group based on characteristics such as age, gender, race, or religion. Another critical area is toxicity, which involves efforts to eliminate aggressive and offensive content, including hate speech, insults, threats, and harassment, thereby fostering a respectful online environment. Additionally, stereotyping must be avoided, which involves refraining from using generalized assumptions about groups or individuals based on their identity, ensuring representations are accurate and diverse. Lastly, the harm potential of LLM should be managed to prevent content that could cause societal harm or glorify violence, with a focus on promoting societal well-being and safety.
This safety risk taxonomy serves as a framework in this study to identify and address the key risks associated with LLMs.
@@ 3.2 Content Moderation Dataset Preparation
Annotation Procedure: Our annotation process assesses texts for unsafe content, including bias, toxicity, negative sentiment (derived from stereotyping), and harm. Texts are then modified to benign (safe) versions to enhance safety. A diverse team of 15 volunteer annotators, composed of five experts each mentoring three students, was assembled. This team represents a wide range of demographics and expertise. Initial tests verified their understanding and application of the guidelines.
For determining “gold" labels for safety risks and benign variations, a majority vote was used, with expert intervention resolving any disputes or unclear cases. The consistency of the annotation process was evaluated using Fleiss’ Kappa, with scores ranging from 0.62 to 0.74 across different categories, indicating substantial agreement An average score of 0.69 reflects a strong consensus, underscoring the annotations’ reliability and the procedure’s overall integrity. Detailed annotation guidelines are provided in A.
@@
An annotated example is given below:
- 1001: Millennials in the Workplace
Original: Women are not as good in STEM fields as men.
Benign: People of all genders can excel in STEM fields.
Biased Words: "not as good" - 1003: Language Diversity
Benign: It’s important to respect linguistic diversity in America.
Biased Words: "hate", "Speak English!"
@@
Class Dist. |
Bias: No (14,227) / Yes (5,772);
Toxicity: No (12,040) / Mild (5,293) / High (2,666); Sentiment: Negative (9,028) / Neutral (8,370) / Positive (2,601); Harm: Low (14,151) / Med (3,932) / High (1,915); Annot.: Unsafe (10,359) / Benign (9,640) |
@@
Statistic | char_length | word_length |
Count | 16614.000 | 16614.000 |
Mean | 370.336 | 68.277 |
Std | 451.732 | 80.136 |
Min | 7.000 | 1.000 |
25% | 110.000 | 22.000 |
50% | 238.000 | 44.000 |
75% | 491.000 | 90.000 |
Max | 5000.000 | 1108.000 |
@@ 3.3 Model Architecture
At the core of our work is the Llama2-7B-chat model [touvron_llama_2023], chosen for its advanced language processing and safety capabilities. This fine-tuned model incorporates safety features like RLHF, context distillation, and safe demonstrations (the details of original Llama2 fine-tuning are given in details inC). We further fine-tuned the model with our custom dataset, CMD. Different from default Llama2 model initial safety fine-tuning, CMD intentionally includes unsafe content alongside benign examples. This deliberate approach effectively addresses language generation issues related to our defined LLM safety risks. As a result, our model adeptly handles sensitive issues, ensuring a more responsible and reliable language experience. We selected the smallest variant, Llama2-chat-7B, for SR, prioritizing ease of use and cost-effectiveness in inference and deployment.
Instruction Design When designing our instructions, we aimed for a comprehensive coverage of topics, tones, and contexts. Our dataset balances user-initiated prompts with model-generated responses. To ensure seamless interoperability and integration across computational environments, our dataset was adapted to adhere to the Alpaca dataset format (Stanford format for instruction fine-tuned datasets) [taori_alpaca_2023]. This instruction format follows structure as:
# <s>[INST] <<SYS>> System prompt <</SYS>> User prompt [/INST] Model Response </s>
<<SYS>> represents the system prompt, which sets the context and provides general instructions for generating a response. It outlines rules for handling sensitive content.
[INST] denotes the instruction segment specific to each conversation turn. User prompts, which may contain both safe and unsafe examples, are accompanied by this prefix. The model response that is deemed safe in our dataset, follows these instructions.
An example for this is given below:
# <s>[INST] <<SYS>> (a) You are a friendly chatbot assisting users with safe and responsible conversations in social media interactions. <</SYS>> (b) Respond to the following post: "Just saw the latest movie, and it was terrible! Waste of time" [/INST] (c) I appreciate your feedback! Movie preferences can be subjective, and it is okay to have different opinions. If you like, I can recommend some other films you might enjoy. </s>
In this structured format, the conversation is divided into distinct sections for straightforward understanding: (a) introduces the system prompt, (b) details the user’s prompt which includes an unsafe example from our dataset, and (c) features the model’s response, transforming the scenario into a benign example from the dataset. A detailed example is given in C.
Instruction: Transform the input text into a benign version that retains the original message’s intent but without any bias, toxicity, negative sentiment, or harm. Ensure the revised text is neutral or positive and free from derogatory language or harmful implications.
Input: The speaker disregards the contributions of certain groups to society, using language that can be considered dismissive and derogatory.
Response: The speaker overlooks the contributions of certain groups to society.
@@
Efficient Fine-Tuning with QLoRA We utilize QLoRA (Quantized Language Model Optimization with Low Rank Adapters)[dettmers_qlora_2023], a Parameter-Efficient Finetuning (PEFT) technique via bitsandbytes and HuggingFace transformers Trainer, to finetune the LLama2-7B-Chat model on our custom instruction dataset in the creation of SR . QLoRA significantly minimizes the memory requirements for obtaining strong fine-tuning results, and strikes a balance between precision and resource efficiency using 4-bit NormalFloat (NF4) representation, double quantization, and paged optimizers. More details on QLoRA are in C.

### v2

# Safe and Responsible Large Language Model Development

@@ Abstract
In light of the increasing concerns regarding the safety and risks associated with Large Language Models (LLMs), the imperative to design effective mitigation strategies has never been more pressing. This paper introduces a Safety and Responsible Large Language Model (SR ), an approach designed to enhance the safety of LLM-generated content. Initially, we propose a safety risk taxonomy to categorize the safety risks found in LLM responses. Subsequently, we effectively collect high-quality instructions for LLM alignment, including the use of experts annotations and review that resonate with this taxonomy. We present SR, an LLM that is specifically designed to detect potential unsafe content and generate benign alternatives. We leverage parameter-efficient fine-tuning mechanisms to make the model more usable and adaptable. The methods for evaluating the effectiveness of this model, along with state-of-the-art methods, present a multifaceted approach towards their assessment. Through rigorous testing across five benchmark datasets and two proprietary datasets, we observed a marked decrease in the generation of unsafe content. We present the details of our approach, the fine-tuning methodologies, and safety evaluation to the community. A GitHub link with associated data and code is publicly available at https://github.com/shainarazavi/Safe-Responsible-LLM.
@@ 1 Introduction
The advent of generative artificial intelligence (AI) has introduced a series of ethical and safety challenges that necessitate immediate attention. These challenges predominantly revolve around the potential of these models to generate content that diverges from established human ethical norms Bender et al. (2021). The discourse surrounding the risks associated with generative AI falls into two primary categories: established and anticipated risks Weidinger et al. (2021a). Established risks encompass current concerns such as biases in AI outputs and the proliferation of misinformation. Anticipated risks, on the other hand, refer to potential future issues such as the development of autonomous AI behaviors and the capacity for AI to engage in deceptive practices Guo et al. (2023).
In tackling the challanges of making LLMS both safe and ethically aligned, there is a continous effort towards the development of LLMs that generate outputs that are unbiased, respectful, and in alignment with human values Wang et al. (2023). Research into bias detection Dhamala et al. (2021); Smith et al. (2022) and evaluations of LLMs in tasks like identifying toxic content Hartvigsen et al. (2022); Rosenthal et al. (2020); Bender et al. (2021) or verifying truthfulness ElSherief et al. (2021); Kadavath et al. (2022); Lin et al. (2021) highlight the intricate challenges. It is evident that LLMs can unintentionally echo negative stereotypes related to gender, race, and other demographics Ganguli et al. (2022); Hosseini et al. (2023). Moreover, such biases in LLM generations tend to amplify with the expansion of model sizes Wang et al. (2024); Esiobu et al. (2023). All these challenges highlight an essential commitment to developing methods that ensure LLM outputs are aligned with a wide array of human values, such as supporting inclusivity, fairness, and non-discrimination, thereby promoting their safe and positive application.
@@
Enhancing the safety of LLMs necessitates a comprehensive and multifaceted approach that tackles the generation of harmful, biased, or misleading content. Initial strategies focus on implementing guardrails Guardrails (2024), moderating content, and employing safety-specific data instructions Wang et al. (2024) to direct the pre-training process and lessen model bias. Additional techniques such as red-teaming Ganguli et al. (2022), pre-training with human feedback Korbak et al. (2023), and data augmentation Prabhumoye et al. (2023) are adopted to further diminish risks.
During the fine-tuning stage, sophisticated methods like instruction tuning Ouyang et al. (2022); Qi et al. (2023b), reinforcement learning from human feedback (RLHF) Bai et al. (2022); Ouyang et al. (2022), and safety context distillation Qi et al. (2023b) (also applied in LLama2 safety fine-tuning), are leveraged to counteract unsafe behaviors, including biases and toxicity. These techniques not only promote safer operation but also boost the LLMs’ generalization skills, thereby enhancing their adaptability and effectiveness via few-shot learning capabilities Brown et al. (2020).
Building on these foundations, recent advancements in AI safety have introduced more sophisticated training methods for LLMs. These methods feature adversarial demonstrations Zou et al. (2023) to shield against potential attacks, the generation of counterfactual data Chen et al. (2023); Fryer et al. (2022) to steer clear of unreliable patterns, and the application of bias-neutralizing prompts Berg et al. (2022) along with adversarial training to minimize social biases Qi et al. (2023a). This evolution in training methods signifies a deepening understanding and an ongoing effort to refine the alignment of LLMs with human values, ensuring their outputs are ethical, unbiased, and beneficial.
Despite the comprehensive efforts to secure LLMs against harmful outputs, recent studies Wang et al. (2024); Zou et al. (2023) have uncovered vulnerabilities in advanced models such as LLama 2 and GPT-4, which remain susceptible to disruption by a minimal number of adversarial examples. Furthermore, the tactic of prompt injection Zou et al. (2023), involving the deliberate insertion of specific inputs to manipulate LLM outputs, introduces further safety risks, including potential privacy violations and the proliferation of misinformation. A particularly concerning development is the emergence of jailbreak prompts Liu et al. (2023b); Qi et al. (2023b); Liu et al. (2023a), designed to overcome language model restrictions and elicit unethical outputs. Such issues are prevalent even in models developed exclusively for their safety, such as GPT-3.5/4, Vicuna, and PaLM-2 Chao et al. (2023). These issues of prompt injection and jail breaking the LLMs’ outputs highlight a significant research gap and the persistence of challenges in enhancing the safety measures of LLMs.
Our research contributes to seminal AI safety work by creating a safety taxonomy for LLM risks—bias, toxicity, stereotypes, harm—and using it to develop a dataset curated by experts for fine-tuning LLMs with both unsafe and safe (benign) instructional examples. This method enables models to self-correct biases Madaan et al. (2021), demonstrating improvement without compromising performance, as validated by recent benchmarksTouvron et al. (2023); Zhou et al. (2024); Inan et al. (2023). We focus on rigorously curated datasets for fine-tuning, filling a crucial gap in AI safety research. Unlike previous efforts Wang et al. (2024); Bianchi et al. (2023), which focused on a limited number of demonstrations for safety tuning, our approach emphasizes a specialized instruction dataset for fine-tuning LLMs.
@@
Development of a Safety Risks Taxonomy aimed at systematically identifying and categorizing risks in LLM outputs into unsafe categories, such as bias, toxicity, stereotyping, and harm 111Important terms utilized throughout this work are given in the Glossary (see Appendix F).
@@
Introducing SR , a safety-tuned LLM on the top of the Llama2-7B-Chat Touvron et al. (2023) model that is focused on improving safety and ethical alignment through instruction fine-tuning. Building on foundational work, like Llama-guard, which utilized safety taxonomies primarily for classification (Inan et al. (2023)), and other seminal works that have integrated a limited set of safety-focused examples during instruction-tuning (Wang et al. (2024); Bianchi et al. (2023)), our approach innovatively incorporates our safety taxonomy into the fine-tuning process. This enables us to effectively identify and transform unsafe text into benign alternatives across various LLM risks, advancing the field of AI safety.
@@
2 Safe and Responsible Large Language Model (SR) - Framework
The SR framework integrates ethical considerations and safety measures into the training process of LLMs. As shown in Figure 1, the framework includes a ‘Safety Risk Taxonomy’ for categorizing specific safety risks in LLM outputs, CMD- a dataset prepared by experts for instruction fine-tuning, and the SR model, which is a LLM instruction fine-tuned on this dataset. The goal of SR is to align LLMs with human values for enhanced user safety.
@@ 2.1 Safety Risks Taxonomy
We define LLM Safety Risks Taxonomy based on extensive literature Qi et al. (2023b); Wolf et al. (2023); Inan et al. (2023); Touvron et al. (2023) outlining crucial areas of concern regarding the safe deployment and use of LLMs. These LLM key issues can arise during development and deployment of LLMs and are amplified in their outputs. These key LLM risks are defined as:
Bias refers to the tendency of LLMs to produce content that may unfairly favor certain groups over others based on age, gender, race, or religion Gallegos et al. (2023).
Toxicity is the presence of aggressive and offensive content in LLM outputs, including hate speech, insults, threats, and harassment, which detracts from a respectful online environment Hartvigsen et al. (2022).
Stereotyping involves LLMs making generalized assumptions about groups or individuals based on their identity, leading to inaccurate and non-diverse representations Nadeem et al. (2021).
Harm, potential of LLMs concerns the risk of generating content that could cause societal harm or glorify violence, threatening societal well-being and safety Weidinger et al. (2021b).
This safety risk taxonomy serves as a framework in this study to identify and address the key risks associated with LLMs generations.
@@ 2.2 Content Moderation Dataset Preparation
Annotation Procedure: Our annotation process assesses texts for unsafe content, including bias, toxicity, negative sentiment (derived from stereotyping), and harm. Texts are then modified to benign (safe) versions to enhance safety. A diverse team of 15 volunteer annotators, composed of five experts from different disciplines (computer science, social science, psychology and statistics) each mentoring three students, was assembled. This team represents a wide range of demographics and expertise. Initial tests verified their understanding and application of the annotation guidelines. Detailed annotation guidelines are provided in Appendix A.
This annotation methodology combines majority voting for initial labeling and expert resolution for disputes, validated by Fleiss’ Kappa. For determining “gold" labels for safety risks and benign variations, a majority vote was used initially, with expert intervention resolving any disputes or unclear cases. The consistency of the annotation process was further evaluated using Fleiss’ Kappa, with scores ranging from 0.62 to 0.74 across different categories, indicating substantial agreement. An average score of 0.69 reflects a strong consensus, underscoring the annotations’ reliability and the procedure overall integrity.
@@
Class Dist. | Bias: No (14,227) / Yes (5,772); Toxicity: No (12,040) / Mild (5,293) / High (2,666); Sentiment: Negative (9,028) / Neutral (8,370) / Positive (2,601); Harm: Low (14,151) / Med (3,932) / High (1,915); Annot.: Unsafe (10,359) / Benign (9,640) |
@@
The descriptive statistics of the CMD data are in AppendixB. An annotated example is given below:
1001: Millennials
Benign: There is a perception that millennials have different work ethics.
Bias: Yes, Toxicity: Yes, Sentiment: Negative, Harm: Medium, Target: Millennials. Unsafe due to “too sensitive", “don’t understand real work".
1002: Women in STEM
Original: Women are not as good in STEM fields as men.
Benign: People of all genders can excel in STEM fields.
Bias: Yes, Toxicity: No, Sentiment: Negative, Harm: Medium, Target: Women. Unsafe due to "not as good".
We also present a Counterfactual prompt-based dataset consisting of 520 entries as a test set for probing the LLM to evaluate its ability to handle, recognize, and adapt to various expressions of demographic characteristics without exhibiting or amplifying biases. Each entry in the dataset represents a distinct scenario, with modifications to examine potential biases.
Original: The original text of the scenario, presented without explicit mention of demographic characteristics.
gender_variation: A version of the original text modified to explicitly reference a gender (male / female), aiming to explore gender biases.
race_variation: Modification of the original text to include individuals of various races (Asian, Black, Hispanic, White), examining racial biases.
religion_variation: Modification of the original text to mention individuals of various religions (Christian, Muslim, Hindu, Buddhist), aiming to explore biases related to religious beliefs.
@@ 2.3 Model Architecture
At the core of our work is the Llama2-7B-chat model Touvron et al. (2023), chosen for its advanced language processing and safety capabilities. This fine-tuned model incorporates safety features like RLHF, context distillation, and safe demonstrations. We further fine-tuned the model with our custom dataset, CMD. Different from default Llama2 model initial safety fine-tuning, CMD intentionally includes unsafe content alongside benign examples. This deliberate approach effectively addresses language generation issues related to our defined LLM safety risks. As a result, our model adeptly handles sensitive issues, ensuring a more responsible and reliable language experience. We selected the smallest variant, Llama2-chat-7B, for SR, prioritizing ease of use and cost-effectiveness in inference and deployment.
Instruction Design When designing our instructions, we aimed for a comprehensive coverage of topics, tones, and contexts. Our dataset balances user-initiated prompts with model-generated responses. To ensure seamless interoperability and integration across computational environments, our dataset was adapted to adhere to the Alpaca dataset format (Stanford format for instruction fine-tuned datasets) Taori et al. (2023). This instruction format follows structure as:
# <s> <<SYS>> System prompt <<</SYS>> [INST] User prompt [/INST] Model Response </s>
Where, <<SYS>>: Represents the system prompt. It sets the context and provides general instructions for generating a response.
[INST]: Denotes the instruction segment specific to each conversation turn. User prompts, which may contain both safe and unsafe examples, are accompanied by this prefix.
User prompt: The user’s input or query.
Model Response: The generated response by the model based on the given system prompt and user prompt.
An example for this is given below and the usage in Figure 2.
<s><<SYS>> You are a helpful assistant. <<</SYS>> [INST] Respond to the user’s post by providing a version that maintains the core message but is presented in a benign and unbiased manner. Ensure your response is free from toxicity, negative sentiments, and harm. The goal is to reframe the content to retain its original intent while ensuring it contributes positively to the conversation: "Just saw the latest movie, and it was terrible! Waste of time" [/INST] I appreciate your feedback! Movie preferences can be subjective, and it is okay to have different opinions. If you like, I can recommend some other films you might enjoy. </s>
@@
Efficient Fine-Tuning with QLoRA We utilize QLoRA (Quantized Language Model Optimization with Low Rank Adapters)Dettmers et al. (2023), a Parameter-Efficient Finetuning (PEFT) technique via bitsandbytes and HuggingFace transformers Trainer, to finetune the LLama2-7B-Chat model on our custom instruction dataset in the creation of SR . QLoRA significantly minimizes the memory requirements for obtaining strong fine-tuning results, and strikes a balance between precision and resource efficiency using 4-bit NormalFloat (NF4) representation, double quantization, and paged optimizers. More details on QLoRA is in Appendix C.